# Collective Superradiance: Estimating the Peak Emission Rate and Time

### v2

@@ IV Universal scalings and bounds
In this section, we show that superradiance in free space follows a universal length scaling and answer the question: what is the maximal spatial extent an emitter ensemble can reach while still being superradiant, namely $R_{\mathrm{peak}}\gtrsim N\Gamma$? Furthermore, we find simple expressions for the variance in Eq. 4, which require just two parameters, the average spacing $\bar{a}$ and total emitter number $N$.
In Fig. 5 (a)-(c) we assume ensembles of $N$ emitters with an equal number of emitters in each spatial dimension, resulting in an ensemble side length $L=\bar{a}N^{1/\mathcal{D}}$ and where $\mathcal{D}$ is the ensemble dimension. The average spacing is obtained by initializing perfect lattices with lattice spacings $a$ and randomly disorder the emitter positions according to a normal distribution with standard deviation $\sigma$. For all plots in Fig. 5 we choose $\sigma/a=2$, however in Appendix C we show, that the variance in Eq. 4 remains mostly unaffected for the even larger positional disorder. In (a)-(c) the variance exhibits the same scaling as a function of the ensemble length $L$ by varying the spacing, irrespective of the emitter number. We define the maximum ensemble length as the value where $\mathrm{Var(\{\Gamma_{k}\}/\Gamma)/N}\sim 10^{-3}$, at which point $R_{\mathrm{peak}}\approx N\Gamma$ according to Eq. 3 and Eq. 4. We also find numerically, that the variances can be well approximated by simple expressions which depend only on the spacing and emitter number, denoted as $\mathrm{Var}_{\mathrm{1D}}$, $\mathrm{Var}_{\mathrm{2D}}$ and $\mathrm{Var}_{\mathrm{3D}}$ in Fig. 5. Expressions are shown in Table 2 along with the bounds on the maximum ensemble lengths. In Fig. 5(d) we study the case of an elongated ensemble in 3D, where $L_{x}>L_{y}$ and find that a maximal superradiant enhancement is reached for $L_{y}\approx 0.2\lambda_{0}$ irrespective of the emitter spacing and length $L_{x}$. The dipole orientations in Fig. 5 are $\mathrm{d}=(0,0,1)^{T}$, however, we find qualitatively the same conclusions for different orientations and large $N$.
@@ V Conclusions and outlook
Acknowledgments - R.H. and S.F.Y. acknowledge NSF via PHY-2207972, the CUA PFC PHY-2317134.
@@ Appendix A The function $\boldsymbol{p(g)}$
Table 3: The $g$ values correspond to values of the second-order correlation $g^{(2)}(0)$ at $t=0$ in Eq. 3. Superradiance occurs only in the interval $21<g<$ 2 and the values of $p(g)$ are the corresponding normalized peak emission rates $p(g)=R_{\mathrm{peak}}/(\Gamma gN^{g})$ obtained from master equation simulations using multiple realizations of various quantum emitter ensembles $(N\leq 12)$ and averaging over the values (see Fig. 1(c) in the main text). This table of values can be interpolated to create the continuous function $p(g)$ in the interval $g\in[1,2]$ and in this work a simple linear interpolation is used.
@@
Figure 6: (a) The function $p(g)$ plotted for $g\in[1,2]$ using a linear interpolation of the data in Table 3 with orange dots representing the normalized peak emission rates in the Dicke limit, $R_{\mathrm{peak}}/(\Gamma gN^{g})$, obtained from the quantum master equation. For $g\in[1,1.75]$ the function $p(g)$ can be approximated by $e^{-2(g-1)}$. (b) A close up for $g\in[1.6,2]$ shows a linear dependence approximated by $p(g)=-0.5g+1.1$.
@@
$p(g)=\begin{cases}\approx\mathrm{e}^{-2(g-1)}&\text{if $1\leq g\lesssim 1.75$}\\ \approx-\frac{1}{2}g+\frac{11}{10}&\text{if $1.75\lesssim g\leq 2$}\end{cases}$ (9)
Since the function $p(g)$ becomes linear in $g$ for $g\gtrsim 1.75$ or in the Dicke limit for $N\geq 8$, it is possible to approximate the peak emission rate for Dicke superradiance as
@@ Appendix B Quantum master equation and dipole-dipole interactions
which results in coherent exchange of excitations reducing the peak emission rate at small spacings, where the coherent exchange rate becomes large, as shown in Fig. 10 and Fig. 11. We note, that the second term in Eq. 10 can be written in diagonal form as
@@ Appendix C Positional disorder
Figure 7: Positional disorder for a single-mode waveguide reservoir. $N=100$ emitters are positioned along the waveguide at positions $\{x_{n}\}$ with waveguide wavenumber $k_{0}=2\pi$, randomly disordered according to a Gaussian distribution with standard deviation $\delta$ around $\{x_{n}\}$ and averaged over $100$ samples. The dots represent the averaged values of Eq. 16 and the continuous lines correspond to Eq. 18. We find excellent agreement for all disorder values and note that $\delta/\lambda_{0}=0.25$ corresponds to a completely disordered ensemble as the waveguide couplings in Eq. 15 can now take any value between 00 and $\Gamma$.
We find numerically, that in the presence of positional disorder of the emitters, $x_{n}+\lambda_{0}\epsilon_{n}$, where $\epsilon_{n}$ is a random displacement with standard deviation $\delta$, expression Eq. 17 can be written as
@@
Fig. 7 shows excellent agreement between Eq. 18 and the exact values obtained from Eq. 16 for increasing positional disorder strengths. We notice that the variance decreases with position disorder as opposed to the free-space case shown in the main text. In particular for $\delta=0.25\lambda_{0}n$, where $n=1,2,3...$, the system is maximally disordered, reducing the variance to the minimum value $(N-2)/2$, which still allows for superradiant decay if the condition $N-2>\Gamma^{\prime}/\Gamma$ is fulfilled. Assuming $\Gamma^{\prime}=0$, the peak emission time is then given by $\tau_{peak}\Gamma=\ln(N/2-1)/(N/2+1)$ and $g=3/2-1/N$ converges to $3/2$ for large $N$, which leads to $R_{\mathrm{peak}}/\Gamma\approx 0.55N^{3/2}$ for a completely disordered chain of $N$ two-level emitters coupled to a single-mode waveguide.
@@ Appendix D Influence of the Hamiltonian on superradiant decay
In Fig. 10 the influence of the Hamiltonian dynamics in Eq. 10 on the peak emission properties is plotted for a ring of emitters in free space and a waveguide reservoir in Fig. 11. The emitters are initially fully excited and decay either with or without the Hamiltonian term present in Eq. B. For smaller emitter spacings in Fig. 10 the difference is more pronounced as the coherent dipole-dipole shifts $J_{nm}$ become substantial and dephases the emission process away from the symmetric Dicke states, while for the peak time a difference arises only in the interval $a\in(0.05,0.15)$. For the waveguide, the Hamiltonian influence increases for $k_{0}a\neq 0$, where the shifts become more pronounced as they depend on the sine of the relative phase. How to accurately quantify the influence of the Hamiltonian on the correlated decay and peak emission properties is an open question at this point and has to be investigated further.
@@
Appendix E Partially excited ensemble in the Dicke limit
In this section, we briefly discuss the case of a partially excited ensemble of the form
$|\psi\rangle_{\varphi}=\bigotimes_{n=1}^{N}\Big{(}\sqrt{1-\varphi}|g_{n}\rangle+\sqrt{\varphi}|e_{n}\rangle\Big{)}=\sum_{m=1}^{N}\varphi^{m/2}(1-\varphi)^{(N-m)/2}\frac{(S^{\dagger})^{m}}{m!}\bigotimes_{n=1}^{N}|g_{n}\rangle,$ (19)
but for fully inverted ensembles with identical decay rates $\Gamma_{n}=\Gamma$, it is again given by Eq. 4 from the main text. The total excitation number at $t=0$ is now given by $N_{\alpha}\leq N$ and the peak emission by
$R_{\varphi}(0)=N\varphi^{2}+N^{2}\varphi(1-\varphi)$ (20)

### v4

@@ IV Universal scalings and bounds
In this section, we show that superradiance in free space follows a universal length scaling and answer the question: what is the maximal spatial extent an emitter ensemble can reach while still being superradiant, namely $R_{\mathrm{peak}}\gtrsim N\Gamma$? Furthermore, we find simple expressions for the variance in Eq. (4), which require just two parameters, the average spacing $\bar{a}$ and total emitter number $N$.
In Fig. 5 (a)-(c) we assume ensembles of $N$ emitters with an equal number of emitters in each spatial dimension, resulting in an ensemble side length $L=\bar{a}N^{1/\mathcal{D}}$ and where $\mathcal{D}$ is the ensemble dimension. The average spacing is obtained by initializing perfect lattices with lattice spacings $a$ and randomly disorder the emitter positions according to a normal distribution with standard deviation $\sigma$. For all plots in Fig. 5 we choose $\sigma/a=2$, however in Appendix C we show, that the variance in Eq. (4) remains mostly unaffected for the even larger positional disorder. In (a)-(c) the variance exhibits the same scaling as a function of the ensemble length $L$ by varying the spacing, irrespective of the emitter number. We define the maximum ensemble length as the value where $\mathrm{Var(\{\Gamma_{k}\}/\Gamma)/N}\sim 10^{-3}$, at which point $R_{\mathrm{peak}}\approx N\Gamma$ according to Eq. (3) and Eq. (4). We also find numerically, that the variances can be well approximated by simple expressions which depend only on the spacing and emitter number, denoted as $\mathrm{Var}_{\mathrm{1D}}$, $\mathrm{Var}_{\mathrm{2D}}$ and $\mathrm{Var}_{\mathrm{3D}}$ in Fig. 5. Expressions are shown in Table 2 along with the bounds on the maximum ensemble lengths. In Fig. 5(d) we study the case of an elongated ensemble in 3D, where $L_{x}>L_{y}$ and find that a maximal superradiant enhancement is reached for $L_{y}\approx 0.2\lambda_{0}$ irrespective of the emitter spacing and length $L_{x}$. The dipole orientations in Fig. 5 are $\mathrm{d}=(0,0,1)^{T}$, however, we find qualitatively the same conclusions for different orientations and large $N$.
@@ V Conclusions and outlook
Acknowledgments - S.F.Y. acknowledge NSF via PHY- 2207972, the CUA PFC PHY-2317134, and QuSeC-TAQS OMA-2326787 in addition to AFOSR FA9550-24-1-0311.
@@ Appendix A The function $\boldsymbol{p(g)}$
Table 3: The $g$ values correspond to values of the second-order correlation $g^{(2)}(0)$ at $t=0$ in Eq. (3). Superradiance occurs only in the interval $21<g<$ 2 and the values of $p(g)$ are the corresponding normalized peak emission rates $p(g)=R_{\mathrm{peak}}/(\Gamma gN^{g})$ obtained from master equation simulations using multiple realizations of various quantum emitter ensembles $(N\leq 12)$ and averaging over the values (see Fig. 1(c) in the main text). This table of values can be interpolated to create the continuous function $p(g)$ in the interval $g\in[1,2]$ and in this work a simple linear interpolation is used.
@@
Figure 6: (a) The function $p(g)$ plotted for $g\in[1,2]$ using a linear interpolation of the data in Table 3 with orange dots representing the normalized peak emission rates in the Dicke limit, $R_{\mathrm{peak}}/(\Gamma gN^{g})$, obtained from the quantum master equation. For $g\in[1,1.8]$ the function $p(g)$ can be approximated by $e^{-2(g-1)}$. (b) A close up for $g\in[1.6,2]$ shows a linear dependence approximated by $p(g)=-0.5g+1.1$.
@@
$p(g)=\begin{cases}\approx\mathrm{e}^{-2(g-1)}&\text{if $1\leq g\lesssim 1.8$}\\ \approx-\frac{1}{2}g+\frac{11}{10}&\text{if $1.8\lesssim g\leq 2$}\end{cases}$ (9)
Since the function $p(g)$ becomes linear in $g$ for $g\gtrsim 1.8$ or in the Dicke limit for $N\geq 8$, it is possible to approximate the peak emission rate for Dicke superradiance as
@@ Appendix B Quantum master equation and dipole-dipole interactions
which results in coherent exchange of excitations reducing the peak emission rate at small spacings, where the coherent exchange rate becomes large, as shown in Fig. 10 and Fig. 11. We note, that the second term in Eq. (10) can be written in diagonal form as
@@ Appendix C Positional disorder
Figure 7: Positional disorder for a single-mode waveguide reservoir. $N=100$ emitters are positioned along the waveguide at positions $\{x_{n}\}$ with waveguide wavenumber $k_{0}=2\pi$, randomly disordered according to a Gaussian distribution with standard deviation $\delta$ around $\{x_{n}\}$ and averaged over $100$ samples. The dots represent the averaged values of Eq. (16) and the continuous lines correspond to Eq. (18). We find excellent agreement for all disorder values and note that $\delta/\lambda_{0}=0.25$ corresponds to a completely disordered ensemble as the waveguide couplings in Eq. (15) can now take any value between 00 and $\Gamma$.
We find numerically, that in the presence of positional disorder of the emitters, $x_{n}+\lambda_{0}\epsilon_{n}$, where $\epsilon_{n}$ is a random displacement with standard deviation $\delta$, expression Eq. (17) can be written as
@@
Fig. 7 shows excellent agreement between Eq. (18) and the exact values obtained from Eq. (16) for increasing positional disorder strengths. We notice that the variance decreases with position disorder as opposed to the free-space case shown in the main text. In particular for $\delta=0.25\lambda_{0}n$, where $n=1,2,3...$, the system is maximally disordered, reducing the variance to the minimum value $(N-2)/2$, which still allows for superradiant decay if the condition $N-2>\Gamma^{\prime}/\Gamma$ is fulfilled. Assuming $\Gamma^{\prime}=0$, the peak emission time is then given by $\tau_{peak}\Gamma=\ln(N/2-1)/(N/2+1)$ and $g=3/2-1/N$ converges to $3/2$ for large $N$, which leads to $R_{\mathrm{peak}}/\Gamma\approx 0.55N^{3/2}$ for a completely disordered chain of $N$ two-level emitters coupled to a single-mode waveguide.
@@ Appendix D Influence of the Hamiltonian on superradiant decay
In Fig. 10 the influence of the Hamiltonian dynamics in Eq. (10) on the peak emission properties is plotted for a ring of emitters in free space and a waveguide reservoir in Fig. 11. The emitters are initially fully excited and decay either with or without the Hamiltonian term present in Eq. (B). For smaller emitter spacings in Fig. 10 the difference is more pronounced as the coherent dipole-dipole shifts $J_{nm}$ become substantial and dephases the emission process away from the symmetric Dicke states, while for the peak time a difference arises only in the interval $a\in(0.05,0.15)$. For the waveguide, the Hamiltonian influence increases for $k_{0}a\neq 0$, where the shifts become more pronounced as they depend on the sine of the relative phase. How to accurately quantify the influence of the Hamiltonian on the correlated decay and peak emission properties is an open question at this point and has to be investigated further.